# Safe Reinforcement Learning on the Constraint Manifold: Theory and Applications

### v1

@@ IV-B Augmented State Dynamics
where $\alpha_{i}:[0,\mu_{\eta})\rightarrow[0,+\infty)$ is a class $\mathcal{K}$ function222class $\mathcal{K}$ function: (1) continuous; (2)strictly increasing; (3)$\alpha(0)=0$ being locally Lipschitz continuous. $\mu_{\eta}$ is a sufficient bigger value that bounds the maximum constraint violation. ${\bm{u}}_{\mu}\in[{\bm{u}}_{\mu}^{-},{\bm{u}}_{\mu}^{+}]\subset\mathbb{R}^{K}$ denotes the vitual control input and ${\bm{u}}_{\mu}^{-}<{\bm{0}}<{\bm{u}}_{\mu}^{+}$. Since the dynamics of each dimension are independent, we develop the following lemma for the 1-dimensional system without loss of generality.
@@
Consider the dynamics $\dot{\mu}=\alpha(\mu){u_{\mu}}$ with $\alpha$ being of class $\mathcal{K}$ and locally Lipschitz for all $\mu\in[0,\mu_{\eta})$. For every initial state $\mu(0)>0$, there exists $\epsilon>0$ such that $\mu(t)\geq\epsilon$, $\forall t\geq 0$, $\forall u_{\mu}\in[u^{-}_{\mu},u^{+}_{\mu}],\,u_{\mu}^{-}<0<u_{\mu}^{+}$.
@@
Lemma 1 shows that the slack variable with the dynamic (8) will not reach zero in finite time if the initial value is not zero.
@@ IV-C Safe Controller on the Tangent Space of the Constraint Manifold
with a constant $\lambda>0$. The first term on the Right Hand Side (RHS) is the Drift Compensation Term, which corrects the drift caused by the system. The middle term is the Contraction Term that retracts the state to the manifold. When $k(\bm{s})<0$, there always exists $\bm{\mu}\in[0,\infty)^{K}$ ensuring $c(\bm{s},\bm{\mu})={\bm{0}}$, so the contraction term will be zero. The last term is the Tangential Term, which generates the vector field tangent to the Constraint Manifold. The overall algorithm is illustrated in Alg. 1. The ATACOM controller can be treated as a generalized solution of the linear system
@@
The tangent space basis ${\bm{B}}_{u}$ can be solved by various approaches, such as QR/SVD decomposition. We will discuss computing a smooth varying basis in Section 6. Next, we will prove that ATACOM enforces the safety constraints under mild assumptions.
Algorithm 1 ATACOM, Input: $\bm{s}$, ${\bm{u}}$triangleright At each step
1:Determine the slack variable $\bm{\mu}=\max(-k(\bm{s}),\mathrm{tol})$
2:Compute the Jacobian ${\bm{J}}_{G},{\bm{J}}_{u}$ and the drift ${\bm{\psi}}$
3:Compute the tangent space basis ${\bm{B}}_{u}$triangleright Alg. 2
4:Compute the constraint value $\bm{c}$
5:Obtain the control output ${\bm{u}}_{s}$triangleright Eq. (12)
@@
Assumption 3 ensures that there exists a non-empty safe set, i.e., $\mathcal{C}\neq\emptyset$. Then, we make another assumption to ensure the set $\mathcal{M}$ is a manifold. From the Constant-Rank Level Set Theorem, we know that the 0-level set $\mathcal{M}$ is a submanifold if the rank of ${\bm{J}}_{u}$ is full row rank. In order to make proper mathematical statements, we first explain some of the notations.
@@
Refer to captionRefer to caption
Figure 4: Illustration of the Constraint Manifolds, Singular Set and the Region of Attraction in Example 3. The thick blue lines depict the constraint manifold, and the red points are singular points. The contraction term shrinks to zero at the singular point, indicating a saddle point of the Lyapunov function $V$. Therefore, we can construct the region of attraction $\Omega_{\eta}$ by excluding the singular set, as the blue shaded area shown in the figure.
@@
In order to present the safety theorem and proof, we introduce the Singular Set, which is used to define the (safe) region of contraction.
@@
where ${\bm{\Sigma}}_{1}$ is the $r\times r$ diagonal matrix whose diagonal entries are the positive singular values of ${\bm{X}}$. The pseudo-inverse of ${\bm{X}}$ is
@@
and
${\bm{X}}{\bm{X}}^{\intercal}={\bm{U}}_{1}{\bm{\Sigma}}_{1}^{2}{\bm{U}}_{1}^{\intercal}$
Since, by hypothesis, ${\bm{x}}^{\intercal}{\bm{X}}{\bm{X}}^{\dagger}{\bm{x}}=0$, we have ${\bm{x}}^{\intercal}{\bm{U}}_{1}{\bm{U}}_{1}^{\intercal}{\bm{x}}=0$. Let ${\bm{y}}={\bm{U}}_{1}^{\intercal}{\bm{x}}$, from the previous equation, we have ${\bm{y}}^{\intercal}{\bm{y}}=0$. This implies ${\bm{y}}={\bm{0}}$, as the dot product of a vector by itself is zero only if the vector is the null vector. We can now write ${\bm{x}}^{\intercal}{\bm{X}}{\bm{X}}^{\intercal}{\bm{x}}={\bm{x}}^{\intercal}{\bm{U}}_{1}{\bm{\Sigma}}_{1}^{2}{\bm{U}}_{1}^{\intercal}{\bm{x}}={\bm{y}}^{\intercal}{\bm{\Sigma}}_{1}^{2}{\bm{y}}=0$. Therefore, we can conclude that ${\bm{X}}^{\intercal}{\bm{x}}={\bm{0}}$. ∎
@@
By definition, we know that ${\bm{J}}_{u}{\bm{B}}_{u}={\bm{0}}$. If Eq. (11) holds, we have ${\bm{\psi}}-{\bm{J}}_{u}{\bm{J}}_{u}^{\dagger}{\bm{\psi}}={\bm{0}}$.
We can simplify the RHS of $\dot{V}$ and get
@@
Due to the negative semi-definiteness of (14), the compact set $\Omega_{\eta}$ is positively invariant. We can find the set $\mathcal{E}\coloneqq\{(\bm{s},\bm{\mu})\in\Omega_{\eta}:\dot{V}=0\}$. From Lemma 2, we have the equivalent sets
@@ IV-D Safety with Stochastic Policy in Reinforcement Learning
Theorem 3 has shown the safety for a Lipschitz continuous controller. However, in reinforcement learning, the control inputs are drawn from a stochastic policy $\pi$ at each time step. The control input is no longer Lipschitz continuous. To deal with the reinforcement learning setting, we treat the system with stochastic policy as a switched system. The system switches to a new one at each time step since a new action is drawn from the stochastic policy.
Many prior works have studied Lasalle’s Invariance principle for switched systems [67, 68, 69]. We present one most relevant here:
@@
Let $\mathcal{M}$ be the union of all the compact weakly invariant sets which are contained in $\mathcal{E}\cap\Omega_{\eta}$. Then every solution $\varphi(t)$ has a nonvanishing dwell time such that $\varphi(0)\in\Omega_{\eta}$ is attracted by $\mathcal{M}$.
@@ V-A Slack Variable Dynamics
Different types of dynamics of the slack variable $\bm{A}$ will affect how the tangent space is deformed. We can use a simple example to understand how the tangent space will be deformed.
@@
Consider a simple constraint $k(s)\coloneqq s\leq 0$ whose Jacobian $J_{k}=1$. Let the dynamics be defined as $\dot{s}=u_{s}$. We get the ${\bm{J}}_{u}=\begin{bmatrix}1&\alpha(\mu)\end{bmatrix}$. We can compute the basis of the tangent space as:
${\bm{B}}_{u}=\frac{1}{\sqrt{1+\alpha^{2}(\mu)}}\begin{bmatrix}-\alpha(\mu)\\ 1\end{bmatrix}$
@@
When designing the slack $\alpha(\mu)$ dynamics, we follow a straightforward principle: The tangential basis should align with the basis of the original system as much as possible when the state is safe and far from the boundary $\partial\mathcal{C}$. Two examples of $\alpha(\mu)$ are
@@
Figure 5 compares a different type of slack dynamics function with different hyperparameters $\beta$. The upper part shows the trajectory between Linear and Exponential dynamics. The lower part shows different hyperparameters $\beta$ for Exponential Slack. A stiff slack dynamics function will deform the action space more aggressively, while a soft slack dynamics function will lead to more conservative behavior. In general, it would be desirable to have as little deformation of the action space as possible to avoid performance loss. However, a stiff slack dynamics function will pose a numerical stability issue as the Jacobian matrix ${\bm{J}}_{u}$ will be ill-conditioned. In addition, a stiff slack dynamics function also requires a higher control frequency as the tangent basis will change more drastically.
@@ V-B Continuously Varying Tangent Space Basis
Figure 6: Comparison of Nonsmooth (N) and Smooth (S) basis using the Linear (LIN) and the Exponential (EXP) slack dynamics function for two different constraints. Top Row: $-(s_{1}^{2}+s_{2}^{2})+1\leq 0$, Bottom Row: $\cos(4s_{1})+s_{2}^{2}-0.8\leq 0$. (a) 3D manifold with linear slack dynamics, the tangent space bases are obtained from QR decomposition. The tangent space bases do not vary smoothly. (b) The tangent basis project onto the original $S_{1}-S_{2}$ space. (c) Projected smooth tangent space bases with linear slack dynamics computed by Alg. 2. The projected tangent space is not orthogonal in the projected space. (d) Projected smooth tangent space bases with exponential slack dynamics. The tangent space bases are less deformed when the state is away from the boundary. (e) The smooth tangent space bases in $S_{1}-S_{2}-\mu$ space.
@@
1:Input: ${\bm{J}}({\bm{x}})$, ${\bm{T}}$
2:Compute the basis matrix ${\bm{B}}({\bm{x}})$ of $\mathrm{ker}{\bm{J}}({\bm{x}})$
3:Compute ${\bm{U}}_{0}\coloneqq({\bm{B}}({\bm{x}}))^{\intercal}{\bm{T}}$
4:Compute the Singular Value Decomposition ${\bm{U}}_{0}={\bm{A}}{\bm{\Sigma}}{\bm{B}}^{\intercal}$
5:Obtain the solution ${\bm{Q}}\coloneqq{\bm{A}}{\bm{B}}^{\intercal}$
6:Form the smooth varying bases ${\bm{B}}^{\prime}({\bm{x}})\coloneqq{\bm{B}}({\bm{x}}){\bm{Q}}$
Let ${\bm{B}}_{u}\in\mathbb{R}^{N\times U}$ be an orthonormal matrix of which columns span the kernel of ${\bm{J}}_{u}\in\mathbb{R}^{S\times N}$ at $(\bm{s},\bm{\mu})\in\mathbb{R}^{N}$. The matrix ${\bm{B}}_{u}$ is not expected to depend continuously on $(\bm{s},\bm{\mu})$ and can be obtained, for example, by QR or SVD decomposition. For any orthogonal matrix ${\bm{Q}}\in\mathbb{R}^{U\times U}$, the matrix ${\bm{B}}_{u}{\bm{Q}}$ is another orthonormal basis of the kernel of ${\bm{J}}_{u}$. The objective is to construct $Q:\mathcal{D}\rightarrow\mathbb{R}^{U\times U}$ such that the ”rotated” bases ${\bm{B}}_{u}{\bm{Q}}$ depend continuously on $(\bm{s},\bm{\mu})$. Rheinboldt [73] suggest formulating an Orthogonal Procrustes Problem as
@@
Consider a velocity-controlled dynamical system $\dot{\bm{s}}={\bm{u}}_{s}$. We compare the tangent bases for two different constraints, $-(s_{1}^{2}+s_{2}^{2})+1\leq 0$ and $\cos(4s_{1})+s_{2}^{2}-0.8\leq 0$. The first constraint constructed a connected manifold, while the second built a disconnected manifold. The constraint manifolds and their tangent space bases obtained from different approaches are shown in Fig. 6. We can observe that the smoothed tangent space bases with exponential slack dynamics are less deformed and continuously varying compared to the ones obtained from QR decomposition with linear slack dynamics.

### v2

@@ IV-B Augmented State Dynamics
where $\alpha_{i}:[0,\mu_{\eta})\rightarrow[0,+\infty)$ is a class $\mathcal{K}$ function222class $\mathcal{K}$ function: (1) continuous; (2)strictly increasing; (3)$\alpha(0)=0$ being locally Lipschitz continuous. $\mu_{\eta}$ is a sufficient big value bounding the maximum constraint violation. ${\bm{u}}_{\mu}\in[{\bm{u}}_{\mu}^{-},{\bm{u}}_{\mu}^{+}]\subset\mathbb{R}^{K}$ denotes the vitual control input and ${\bm{u}}_{\mu}^{-}<{\bm{0}}<{\bm{u}}_{\mu}^{+}$. Since the dynamics of each dimension are independent, we develop the following lemma for the 1-dimensional system without loss of generality.
@@
Consider the dynamics $\dot{\mu}=\alpha(\mu){u_{\mu}}$ with $\alpha$ being of class $\mathcal{K}$ and locally Lipschitz for all $\mu\in[0,\mu_{\eta})$, where $\mu_{\eta}<\infty$ is a sufficient big value that upper bounds $\mu$. For every initial state $\mu(0)>0$, there exists $\epsilon>0$ such that $\mu(t)\geq\epsilon$, $\forall t\geq 0$, $\forall u_{\mu}\in[u^{-}_{\mu},u^{+}_{\mu}],\,u_{\mu}^{-}<0<u_{\mu}^{+}$.
@@
Lemma 1 shows that the slack variable with the dynamics basis of the (8) will not reach zero in finite time if the initial value is not zero.
@@ IV-C Safe Controller on the Tangent Space of the Constraint Manifold
Algorithm 1 Safe Reinforcement Learning with ATACOM
1:Initialize: policy $\pi$, replay buffer $\mathcal{D}$, number of steps $N$
2:for $k\leftarrow 1$ to $N$  do
3:     Sample action ${\bm{u}}\sim\pi(\cdot|\bm{s})$.
4:     Obtain safe control ${\bm{u}}_{s}\leftarrow\text{ATACOM}({\bm{s}},{\bm{u}})$. $\triangleright$ Alg. 2
5:     Execute ${\bm{u}}_{s}$ and obtain $\bm{s}^{\prime}$ and reward $r$.
6:     Update replay buffer $\mathcal{D}\leftarrow(\bm{s},{\bm{u}},\bm{s}^{\prime},r)$.
7:     Update RL agent using the replay buffer $\mathcal{D}$.
8:end for
@@
with a constant $\lambda>0$. The first term on the Right Hand Side (RHS) is the Drift Compensation Term, which corrects the drift caused by the system. The middle term is the Contraction Term that retracts the state to the manifold. When $k(\bm{s})<0$, there always exists $\bm{\mu}\in[0,\infty)^{K}$ ensuring $c(\bm{s},\bm{\mu})={\bm{0}}$, so the contraction term will be zero. The last term is the Tangential Term, which generates the vector field tangent to the Constraint Manifold. The ATACOM controller can be treated as a generalized solution of the linear system
@@
The tangent space basis ${\bm{B}}_{u}$ can be solved by various approaches, such as QR/SVD decomposition. We will discuss computing a smooth varying basis in Section V-B. Alg. 1 and Alg. 2 illustrate how to integrate ATACOM into the RL framework. Generally speaking, ATACOM constructs a safe action space that morphs all input actions into safe actions. Therefore, the RL agent will learn the policy directly from the safe action space instead of the original one.
Algorithm 2 ATACOM
1:Input: $\bm{s}$, ${\bm{u}}$triangleright At each step
2:Determine the slack variable
3:$\bm{\mu}\leftarrow\max(-k(\bm{s}),\mathrm{tol})$
4:Compute the Jacobians and the drift
5:${\bm{J}}_{G}\leftarrow J_{k}(\bm{s})G(\bm{s})$
6:${\bm{J}}_{u}(\bm{s},\bm{\mu})\leftarrow\begin{bmatrix}{\bm{J}}_{G}(\bm{s})&A(\bm{\mu})\end{bmatrix},$
7:${\bm{\psi}}(\bm{s})\leftarrow J_{k}(\bm{s})f(\bm{s})$
8:Compute the tangent space basis $\triangleright$ Alg. 3
9:${\bm{B}}_{u}\leftarrow\text{SmoothBasis}({\bm{J}}_{u})$
10:Compute the constraint value
11:$c(\bm{s},\bm{\mu})\leftarrow k(\bm{s})+\bm{\mu}$
12:Compute safe control output ${\bm{u}}_{s}$triangleright Eq. (12)
13:Output: ${\bm{u}}_{s}$
Next, we will prove that ATACOM enforces the safety constraints under mild assumptions.
@@
Assumption 3 ensures that there exists a non-empty safe set, i.e., $\mathcal{C}\neq\emptyset$. Then, we make another assumption to ensure the set $\mathcal{M}$ is a manifold. From the Constant-Rank Level Set Theorem, the 0-level set $\mathcal{M}$ is a submanifold if the rank of ${\bm{J}}_{u}$ is full row rank. To make proper mathematical statements, we first explain some of the notations.
@@
Remark 3.
Assumption 4 is a necessary condition to ensure that (7) forms a manifold, as dictated by the constant-rank level set theorem. It is also a sufficient condition to guarantee the existence of a solution to the linear system (11). However, when considering the input constraint ${\bm{u}}_{s}\in\mathcal{U}$, this assumption alone is not sufficient to ensure the existence of a solution.
To present the safety theorem and proof, we introduce the Singular Set, used in defining the (safe) region of contraction.
@@
Figure 3: Illustration of the Constraint Manifolds, Singular Set, and the Region of Attraction in Example 3. The thick blue lines depict the constraint manifold, and the red points are singular points. The contraction term shrinks to zero at the singular point, indicating a saddle point of the Lyapunov function $V$. The region of attraction $\Omega_{\eta}$ excludes the singular set, as the blue shaded area shown in the figure.
@@
where ${\bm{\Sigma}}_{1}$ is the $r\times r$ diagonal matrix whose diagonal entries are the positive singular values of ${\bm{X}}$. The pseudo-inverse is
@@
and ${\bm{X}}{\bm{X}}^{\intercal}={\bm{U}}_{1}{\bm{\Sigma}}_{1}^{2}{\bm{U}}_{1}^{\intercal}$. Since, by hypothesis, ${\bm{x}}^{\intercal}{\bm{X}}{\bm{X}}^{\dagger}{\bm{x}}=0$, we have ${\bm{x}}^{\intercal}{\bm{U}}_{1}{\bm{U}}_{1}^{\intercal}{\bm{x}}=0$. Let ${\bm{y}}={\bm{U}}_{1}^{\intercal}{\bm{x}}$, from the previous equation, we have ${\bm{y}}^{\intercal}{\bm{y}}=0$. This implies ${\bm{y}}={\bm{0}}$, as the dot product of a vector by itself is zero only if the vector is the null vector. We can now write ${\bm{x}}^{\intercal}{\bm{X}}{\bm{X}}^{\intercal}{\bm{x}}={\bm{x}}^{\intercal}{\bm{U}}_{1}{\bm{\Sigma}}_{1}^{2}{\bm{U}}_{1}^{\intercal}{\bm{x}}={\bm{y}}^{\intercal}{\bm{\Sigma}}_{1}^{2}{\bm{y}}=0$. Therefore, we can conclude that ${\bm{X}}^{\intercal}{\bm{x}}={\bm{0}}$. ∎
@@
By definition, we know that ${\bm{J}}_{u}{\bm{B}}_{u}={\bm{0}}$. If Eq. (11) holds, we have ${\bm{\psi}}-{\bm{J}}_{u}{\bm{J}}_{u}^{\dagger}{\bm{\psi}}={\bm{0}}$. We can simplify the RHS and get
@@
Due to the negative semi-definiteness of (14), the compact set $\Omega_{\eta}$ is positively invariant. We can find the set $\mathcal{E}\coloneqq\{(\bm{s},\bm{\mu})\in\Omega_{\eta}:\dot{V}=0\}$. From Lemma 2, we have the equivalent set
@@ IV-D Safety with Stochastic Policy in Reinforcement Learning
Theorem 3 has shown the safety for a Lipschitz continuous controller. However, in reinforcement learning, the control inputs are drawn from a stochastic policy $\pi$ at each time step. The control input is no longer Lipschitz continuous. To deal with the reinforcement learning setting, we treat the system with stochastic policy as a switched system. The system switches to a new one at each time step when a new action is drawn from the stochastic policy. Many prior works have studied Lasalle’s Invariance principle for switched systems [68, 69, 70]. We present the most relevant one here:
@@
Let $\mathcal{M}$ be the union of all compact weakly invariant sets contained in $\mathcal{E}\cap\Omega_{\eta}$. Every solution $\varphi(t)$ has a nonvanishing dwell time such that $\varphi(0)\in\Omega_{\eta}$ is attracted by $\mathcal{M}$.
@@
IV-E Input-to-State Stability with disturbances
In the previous analysis, we assume full knowledge of the dynamics is available. In practice, a perfect dynamic model is not available due to model mismatch and disturbances. In this section, we address this issue by considering the dynamic model with external disturbances as
$\bm{s}=f(\bm{s})+G(\bm{s}){\bm{u}}_{s}+\bm{\epsilon}$ (15)
where $\|\bm{\epsilon}\|<\omega$ is a bounded random variable that captures the model mismatch and external disturbances. In the following theorem, we prove the dynamic system with disturbances (15) is Input-to-State Stable (ISS) [71].
Theorem 5.
Consider the dynamic system (15) and the safety constraint (5). Let $\mathcal{M}$ be the constraint manifold defined in (7) and the dynamics of the slack variable defined in (8). The system with controller (12) is Input-to-State Stable. Every trajectory, starting from $\Omega_{\eta}$, will approach the neighborhood of constraint manifold $\{(\bm{s},\bm{\mu})|\|\bm{c}\|=\|k(\bm{s})+\bm{\mu}\|\leq\eta_{c},\sqrt{2\eta}>\eta_{c}>0\}$, if $\|{\bm{J}}_{k}\|\leq\eta_{J}$, $\lambda\geq\frac{\eta_{J}}{\eta_{c}}\omega$, and (11) holds.
Proof.
Consider the Lyapunov-like function $V=\frac{1}{2}\bm{c}^{\intercal}\bm{c}$. Following a similar derivation from (14), we have $\dot{V}=-\lambda\bm{c}^{\intercal}{\bm{J}}_{u}{\bm{J}}_{u}^{\dagger}\bm{c}+\bm{c}^{\intercal}{\bm{J}}_{k}\bm{\epsilon}$. From Assumption 4, ${\bm{J}}_{u}$ is full row rank and ${\bm{J}}_{u}{\bm{J}}_{u}^{\dagger}=\mathbb{I}$. We have
$\dot{V}$ $\leq-\lambda\|\bm{c}\|^{2}+\|\bm{c}\|\|{\bm{J}}_{u}\|\|\bm{\varepsilon}\|$
The system is ISS following [71] [Theorem 1]. When $\sqrt{2\eta}>\|\bm{c}\|\geq\eta_{c}$
$\dot{V}$ $\leq-\frac{\omega\eta_{J}}{\eta_{c}}\|\bm{c}\|^{2}+\omega\eta_{J}\|\bm{c}\|<0$
The system starting from $(\bm{s},\bm{\mu})\in\Omega_{\eta}$ will approach the neighborhood of the manifold $\{(\bm{s},\bm{\mu})|\|k(\bm{s})+\bm{\mu}\|\leq\eta_{c}\}$. Given that $\|\bm{k}\|\leq\|\bm{c}\|$. The maximum constraint violation will also be bounded by $\eta$. ∎
@@ V-A Slack Variable Dynamics
Different types of dynamics of the slack variable $\bm{A}$ will affect how the tangent space is deformed. We use a simple example to understand how the tangent space is deformed.
@@
Consider a simple constraint $k(s)\coloneqq s\leq 0$ whose Jacobian $J_{k}=1$. Let the dynamics be defined as $\dot{s}=u_{s}$. We get the ${\bm{J}}_{u}=\begin{bmatrix}1&\alpha(\mu)\end{bmatrix}$. We can compute the basis of the tangent space as ${\bm{B}}_{u}=\left[-\alpha(\mu)\quad 1\right]^{\intercal}/\sqrt{1+\alpha^{2}(\mu)}$.
@@
When designing the slack $\alpha(\mu)$ dynamics, we follow a straightforward principle: The tangential basis should align with the basis of the original system as much as possible when the state is safe and far from the boundary $\partial\mathcal{C}$. For example,
@@
Fig. 4 compares a different type of slack dynamics function with different hyperparameters $\beta$. The upper part shows the trajectory between Linear and Exponential dynamics. The lower part shows different hyperparameters $\beta$ for Exponential Slack. A stiff slack dynamics function will deform the action space more aggressively, while a soft slack dynamics function will lead to more conservative behavior. In general, it would be desirable to have as little deformation of the action space as possible to avoid performance loss. However, a stiff slack dynamics function will pose a numerical stability issue as the Jacobian matrix ${\bm{J}}_{u}$ will be ill-conditioned. In addition, a stiff slack dynamics function also requires a higher control frequency as the tangent basis will change more drastically.
@@ V-B Continuously Varying Tangent Space Basis
Figure 5: Comparison of Nonsmooth (N) and Smooth (S) basis using the Linear (LIN) and the Exponential (EXP) slack dynamics function for two different constraints. Top Row: $-(s_{1}^{2}+s_{2}^{2})+1\leq 0$, Bottom Row: $\cos(4s_{1})+s_{2}^{2}-0.8\leq 0$. (a) 3D manifold with linear slack dynamics, the tangent space bases are obtained from QR decomposition. The tangent space bases do not vary smoothly. (b) The tangent basis project onto the original $S_{1}-S_{2}$ space. (c) Projected smooth tangent space bases with linear slack dynamics computed by Alg. 3. The projected tangent space is not orthogonal in the projected space. (d) Projected smooth tangent space bases with exponential slack dynamics. The tangent space bases are less deformed when the state is away from the boundary. (e) The smooth tangent space bases in $S_{1}-S_{2}-\mu$ space.
@@
1:Input: ${\bm{J}}$, Initialize: ${\bm{T}}$
2:Compute the basis matrix using SVD/QR decomposition.
3:${\bm{B}}\leftarrow\mathrm{ker}{\bm{J}}$
4:Compute ${\bm{U}}_{0}\leftarrow{\bm{B}}^{\intercal}{\bm{T}}$.
5:Compute ${\bm{A}},{\bm{\Sigma}},{\bm{B}}^{\intercal}\leftarrow\text{SVD}({\bm{U}}_{0})$
6:Obtain the solution ${\bm{Q}}\leftarrow{\bm{A}}{\bm{B}}^{\intercal}$.
7:Form the smooth varying bases ${\bm{B}}^{\prime}\leftarrow{\bm{B}}{\bm{Q}}$.
8:Output: ${\bm{B}}^{\prime}$
Let ${\bm{B}}_{u}\in\mathbb{R}^{N\times U}$ be an orthonormal matrix of which columns span the kernel of ${\bm{J}}_{u}\in\mathbb{R}^{S\times N}$ at $(\bm{s},\bm{\mu})\in\mathbb{R}^{N}$. The matrix ${\bm{B}}_{u}$ is not expected to depend continuously on $(\bm{s},\bm{\mu})$ and can be obtained, for example, by QR or SVD decomposition. For any orthogonal matrix ${\bm{Q}}\in\mathbb{R}^{U\times U}$, the matrix ${\bm{B}}_{u}{\bm{Q}}$ is another orthonormal basis of $\mathrm{ker}{\bm{J}}_{u}$. The objective is to construct $Q:\mathcal{D}\rightarrow\mathbb{R}^{U\times U}$ such that the ”rotated” bases ${\bm{B}}_{u}{\bm{Q}}$ depend continuously on $(\bm{s},\bm{\mu})$. Rheinboldt [75] suggest formulating an Orthogonal Procrustes Problem as
@@
Consider a velocity-controlled dynamical system $\dot{\bm{s}}={\bm{u}}_{s}$. We compare the tangent bases for two different constraints, $-(s_{1}^{2}+s_{2}^{2})+1\leq 0$ and $\cos(4s_{1})+s_{2}^{2}-0.8\leq 0$. The first constraint constructs a connected manifold, while the second builds a disconnected one. The constraint manifolds and their tangent space bases obtained from different approaches are shown in Fig. 5. We can observe that the smoothed tangent space bases with exponential slack dynamics are less deformed and continuously varying compared to the ones obtained from QR decomposition with linear slack dynamics.